write equation of parabola in vertex form with vertex:(1,4) Focus:(-1,4)
Answer by lwsshak3(11628)

write equation of parabola in vertex form with vertex:(1,4) Focus:(-1,4)
**
Basic form of equation for a parabola that opens leftward: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex.
For given problem:
vertex: (1,4)
axis of symmetry: y=4
p=2 (distance from focus to vertex on the axis of symmetry)
4p=8
Equation: (y-4)^2=-8(x-1)
(y-4)^2=-8x+8
-8x=(y-4)^2-8
x=-(1/8)(y-4)^2+1 (vertex form of equation)